OpenStudy (anonymous):

12z - 3 is greater than or the same as 2z - 15

OpenStudy (anonymous):

\[12z-3\ge2z-15\] ==>\[10z \ge-12\] ==>\[z \ge-1.2\]

OpenStudy (anonymous):

see thats the answer i got but its wrong.. :/ its supposed to be a fraction...

OpenStudy (anonymous):

-6/5

OpenStudy (anonymous):

Thank you!!!!!!!!!!!

OpenStudy (anonymous):

you're always welcome

OpenStudy (anonymous):

hey how'd you change it into fraction?

OpenStudy (anonymous):

seem my last step,i got, \[10z \ge-12\] therefore,==>\[z \ge-12/10\] cutting common factor 2 from top and bottom, ==>\[z \ge-6/5\]

OpenStudy (anonymous):

okay thanks.. :D

OpenStudy (anonymous):

..D